OpenStudy (anonymous):

i have to find the x and y intercept of these questions so can anybody EXPLAIN to me how to do them and not just give me the answers? thanks you to anyone, y=-7x y=1/2x=3 -2y=5x-12

OpenStudy (anonymous):

You ALWAYS need to put it in standard form first. Standard form: y=mx+b So, m= slope b= y-intercept

OpenStudy (anonymous):

and x=variable

OpenStudy (anonymous):

so then it would be y=-7x+b?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

And when finding the x intercept, make y zero And when finding the y intercept make x zero :)

OpenStudy (anonymous):

0=-7+b right? then what would be after that?

OpenStudy (anonymous):

Yes :-) So then add 7 to both sides so you get the variable on one side :)

OpenStudy (anonymous):

Make sure it is positive 7 though!

OpenStudy (anonymous):

7x or just 7?

OpenStudy (anonymous):

Just 7 :)

OpenStudy (anonymous):

7=x+b?

OpenStudy (anonymous):

Yes. so make x zero now to find the y intercept :)

OpenStudy (anonymous):

do i start from the beginning again or just keep going off from the 7=x+b?

OpenStudy (anonymous):

Keep going on that equation, and make x=0., does that make sense? :)

OpenStudy (anonymous):

yes so it would be 7=0+b

OpenStudy (anonymous):

Yep! Now get b by itself to find the y intercept when x is 0 :) That will give you your first ordered pair!

OpenStudy (anonymous):

how do i get b by itself?

OpenStudy (anonymous):

Well actually I messed up :) You already have b by itself, so when x is zero y=7.

OpenStudy (anonymous):

So the first ordered pair would be (0,7)

OpenStudy (anonymous):

ok so then that is the answer for the first one?

OpenStudy (anonymous):

Yes :) Now do you want me to work through the next one with you? :)

OpenStudy (anonymous):

yea because its different because it has the 1/2 and everything

OpenStudy (anonymous):

Lol ok :)

OpenStudy (anonymous):

so i think the first thing i would do is turn 1/2 into 0?

OpenStudy (anonymous):

Yeah, because it would be 1/2 times 0 which equals zero :)

OpenStudy (anonymous):

y=0x=3?

OpenStudy (anonymous):

Yep :-)

OpenStudy (anonymous):

So when x is zero y is 3, right? :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So now you just found another ordered pair! (0,3) Do you understand what the x and y intercept means?

OpenStudy (anonymous):

yes and thank you for helping me with these problems. with the last one though what would i do since the y has -2 with it?

OpenStudy (anonymous):

You are welcome, it is my pleasure! Let me look at the last problem here real quick :)

OpenStudy (anonymous):

Ok so to find the x and y intercepts we do what we just did. Make x zero to find the y intercept and then i will help you with what to do afterwards :)

OpenStudy (anonymous):

-2y=0-12

OpenStudy (anonymous):

Yep :-) Sorry I was doing something real quick. So -2y=-12. Now divide each side by (-2)